# Preserving Angles Improves Feature Distillation of Foundation Models

### v2

@@ 6 Conclusion
This paper introduces CosPress, a feature distillation approach designed to train highly performant student networks from a foundation model teacher with a Vision Transformer [11] architecture. This is achieved by introducing a teacher head, that maps from the higher dimensional latent space of the teacher network into the smaller dimensional space of the student, and training this mapping to preserve the cosine similarity of images within these embedding spaces. CosPress trains a faithful student, that more closely replicates the behaviour of the teacher network in comparison to the Proteus approach [56], where a student head is used to align the student outputs with the teacher.
@@ Acknowledgements
We acknowledge the Traditional Custodians of the unceded land on which the research detailed in this paper was undertaken, the Wurundjeri Woi Wurrung and Bunurong peoples of the Kulin nation, and pay our respects to their Elders past and present. This research was undertaken using the LIEF HPC-GPGPU Facility hosted at the University of Melbourne. This Facility was established with the assistance of LIEF Grant LE170100200. This research was also undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. Evelyn J. Mannix was supported by a Australian Government Research Training Program Scholarship to complete this work.
@@ Appendix A The Johnson–Lindenstrauss Lemma
The Johnson–Lindenstrauss (JL) Lemma [13] states that for a set of points $\mathbf{X}$ in a high dimensional space, there exists a function that can map these points into a lower dimensional space, within error $\epsilon$, where this error depends on the dimension of the target space $m$ and the size of the set of points $|\mathbf{X}|$. In its standard form, the JL lemma states that Euclidean distances are preserved.
Lemma 2 (Johnson–Lindenstrauss; [13]).
@@
Lemma 3 (Norm preservation; [13]).
@@
Again, for target spaces with larger dimension $m$ this Lemma 3 states that it is more likely that a high quality map will be sampled. A similar result also holds for angles [26], which gives
Lemma 4 (Angles; [26]).
@@
The proof of Lemma 4 also relies on $f$ being generated as a random projection as above [26]. The JL transform matrix $\mathbf{M}$ has a further interesting property, as observed in this work, that we refer to as approximate left and right orthogonality.
@@
and introducing Eq. 29 from above obtains
@@
which proves the first statement. It is easy to see the converse inequality is not true, simply by constructing a matrix $\mathbf{M}\in\mathbb{R}^{d\times m}$ with the first $m$ rows equal to the identity. This matrix satisfies that $\mathbf{M}\mathbf{M}^{\top}=\mathbf{I}_{m}$, but $\mathbf{M}^{\top}\mathbf{M}$ will have $d-m$ rows with only zeros in them. This proves the second statement. ∎
@@ Appendix B Further Results
Table S11: Specialist models — generalisability. Comparison of generalisability of specialist models on the cartoon subsets of the CIFAR-10-W benchmark [38]. We report mean per-class accuracy due to dataset imbalances. In-distribution training images (top) and cartoon images (bottom) are included for reference.
Method Arch Teacher Pretraining dataset Linear DeiT
Diff Bin Bai 360 Diff Bin Bai 360
DeiT [40] ViT-Ti/16 RegNetY-16GF ImageNet 65.9 51.0 47.6 48.8 86.9 62.6 56.0 60.1
Table S12: Specialist models — OOD detection. Comparison of performance on the OpenOOD benchmark [48]. The AUC is reported for detecting OOD images.
CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
DeiT [40] ViT-Ti/16 RegNetY-16GF ImageNet 57.01 47.04 58.14 46.19 96.79 98.69 87.45 86.73
The challenging cartoon subsets of the CIFAR-10-W benchmark [38] are used to test generalisation performance on CIFAR-10. Tab. S11 shows that CosPress finetuning leads to improved generalisability for specialist models on CIFAR-10. Under a linear probing evaluation, CosPress finetuning strongly improves generalisability on two of the four datasets, and improves generalisability for all datasets even after DeiT finetuning. Tab. S11 further shows consistent generalisability improvements over the baseline DeiT pretrained model.
DeiT finetuning is observed to significantly improve generalisation performance across all initializations, due to extensive use of both mixup [54] and cutmix [52] in the DeiT training recipe.
The CIFAR-10/100 OpenOOD benchmarks [48] are used to test OOD detection performance. Tab. S12 shows that CosPress finetuning leads to improved performance on the OpenOOD [48] benchmark for specalist models on CIFAR-10/100. Without DeiT finetuning, strong improvements in OOD detection are observed under CosPress finetuning over the ImageNet pretrained baselines. Improvements remain after DeiT finetuning, but are smaller.
DeiT finetuning is also observed to significantly improve OOD detection across all initializations, which is consistent with prior work showing cutmix [52] to improve performance on these benchmarks.
@@
We observe that training with larger teachers in comparison to the student requires longer training runs. To achieve best results efficiently, the pretrained models from a smaller teacher are used as a starting point. Then, the student and teacher heads are first trained with a frozen pretrained student model (allowing the CosPress teacher heads to minimize the student loss) for 30 epochs. Finally, the models are trained for 300 epochs using the same distillation approach as previously. Tab. S13 shows that this improves the performance of the student models, with CosPress seeing larger improvements in accuracy on ImageNet-1K in comparison to Proteus.
However, this results in poorer results in dense image tasks, like semantic segmentation. Tab. S14 shows that the students trained with a larger teacher network have poorer mIoU for a linear probe on the Pascal VOC 2012 dataset. Undertaking longer distillation runs, or continuing training using a larger image resolution potentially might be helpful in these cases.
Method Arch Teacher kNN Linear
Backbone Student head Teacher head Teacher
Proteus ViT-Ti/14 DINOv2 ViT-S/14 73.1 73.5 79.0 76.1
CosPress ViT-Ti/14 DINOv2 ViT-S/14 74.3 78.8 79.0 76.8
Proteus ViT-Ti/14 DINOv2 ViT-B/14 73.4 73.8 82.1 76.9
CosPress ViT-Ti/14 DINOv2 ViT-B/14 75.6 81.9 82.1 77.9
@@
Method Arch Teacher mIoU
Proteus ViT-Ti/14 DINOv2 ViT-S/14 70.5
CosPress ViT-Ti/14 DINOv2 ViT-S/14 71.1
Proteus ViT-Ti/14 DINOv2 ViT-B/14 68.1
CosPress ViT-Ti/14 DINOv2 ViT-B/14 69.7
@@
The OOD detection results for all of the OpenOOD datasets [48] are presented in Tab. S15, Tab. S16 and Tab. S17. The tables report the ROC-AUC for detecting OOD images and the False Positive Rate (FPR) using a 95% threshold for including all in-distribution images. To produce these results, the KNN+ [39] OOD metric is used to measure the performance of the model backbones. For ImageNet-1K, we sample 1% of the dataset (12,812 images) and set $k=10$ to measure distance to OOD samples. For CIFAR-10/100, we sample 100% of the dataset (50,000 images) and set $k=1$.
@@
Table S16: Specialist models — near OOD detection. Comparison of performance on the OpenOOD benchmark [48]. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
@@
Table S17: Specialist models — far OOD detection. Comparison of performance on the OpenOOD benchmark [48]. The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
@@ Appendix C Ablation Studies
Method Arch Teacher $\gamma$ kNN
CosPress ViT-Ti/14 DINOv2 ViT-S/14 - 74.3
CosPress ViT-Ti/14 DINOv2 ViT-S/14 10 74.3
CosPress ViT-Ti/14 DINOv2 ViT-S/14 100 74.2
@@
We consider the performance of weighting the dimensionality reduction loss in Eq. 6, rather than freezing the gradients of $\phi$ in the student loss. This gives an alternative loss function
@@
where $\gamma$ is a weighting factor prioritises the dimensionality reduction loss when it is set to be greater than one. Tab. S18 shows that the approach of weighting or freezing gradients leads to similar results.
@@
Method Arch Teacher Loss Metric kNN
CosPress ViT-Ti/14 DINOv2 ViT-S/14 Cosine distance 74.3
CosPress ViT-Ti/14 DINOv2 ViT-S/14 MSE 74.2
@@
Tab. S19 considers the impact on performance of using different metrics for the student loss $\mathcal{L}_{\text{student}}$ for Eq. 14. It is found that using a cosine distance loss leads to slightly better performance in comparison to a mean squared error loss.
@@
Method Arch Teacher $\bm{\tau}$ kNN
CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\bm{\left[0.01,0.02,0.03,0.04,0.05,0.06,0.07,0.08,0.09,0.10\right]}$ 74.3
CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\left[0.01\right]$ 74.1
CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\left[0.10\right]$ 74.5
CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\left[0.01,0.10\right]$ 74.3
@@
Tab. S19 shows the performance impact of different sets of temperatures $\bm{\tau}$ in the dimensionality reduction loss for Eq. 10. It is found that these values have a small impact on performance, with the best set being $\bm{\tau}=\left[0.10\right]$, which obtains slightly better performance that the set of parameters chosen for the paper.

### v3

@@ 7 Conclusion
This paper introduces CosPress, a feature distillation approach designed to train highly performant student networks from a foundation model teacher with a Vision Transformer (dosovitskiy2020image) architecture, that reproduces their properties in regards to generalisation, robustness and OOD detection. This is achieved by introducing a teacher head, that maps from the higher dimensional latent space of the teacher network into the smaller dimensional space of the student, and training this mapping to preserve the cosine similarity of images within these embedding spaces. CosPress trains a faithful student, that more closely replicates the behaviour of the teacher network in comparison to the Proteus approach (zhang2024accessingvisionfoundationmodels), where a student head is used to align the student outputs with the teacher.
@@ Acknowledgements
We acknowledge the Traditional Custodians of the unceded land on which the research detailed in this paper was undertaken, the Wurundjeri Woi Wurrung and Bunurong peoples of the Kulin nation, and pay our respects to their Elders past and present. This research was undertaken using the LIEF HPC-GPGPU Facility hosted at the University of Melbourne. This Facility was established with the assistance of LIEF Grant LE170100200. This research was also undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. Evelyn J. Mannix was supported by an Australian Government Research Training Program Scholarship to complete this work.
@@ Appendix A The Johnson–Lindenstrauss Lemma
The Johnson–Lindenstrauss (JL) Lemma (freksen2021introduction) states that for a set of points $\mathbf{X}$ in a high dimensional space, there exists a function that can map these points into a lower dimensional space, within error $\epsilon$, where this error depends on the dimension of the target space $m$ and the size of the set of points $|\mathbf{X}|$. In its standard form, the JL lemma states that Euclidean distances are preserved.
Lemma 2 (Johnson–Lindenstrauss; (freksen2021introduction)).
@@
Lemma 3 (Norm preservation; (freksen2021introduction)).
@@
Again, for target spaces with larger dimension $m$ this Lemma 3 states that it is more likely that a high quality map will be sampled. A similar result also holds for angles (magen2007dimensionality), which gives
Lemma 4 (Angles; (magen2007dimensionality)).
@@
The proof of Lemma 4 also relies on $f$ being generated as a random projection as above (magen2007dimensionality). The JL transform matrix $\mathbf{M}$ has a further interesting property, as observed in this work, that we refer to as approximate left and right orthogonality.
@@
and introducing Eq.˜30 from above obtains
@@
which proves the first statement. It is easy to see the converse inequality is not true, simply by constructing a matrix $\mathbf{M}\in\mathbb{R}^{d\times m}$ with the first $m$ rows equal to the identity. This matrix satisfies that $\mathbf{M}\mathbf{M}^{\top}=\mathbf{I}_{m}$, but $\mathbf{M}^{\top}\mathbf{M}$ will have dmd-m rows with only zeros in them. This proves the second statement. ∎
Algorithm 1 Algorithm for feature distillation with CosPress.
Input: Training set $X$, teacher model $T$, student model $S_{\theta}$, teacher head $h_{\phi}$, $N_{E}$ number of epochs, $\text{Aug}(.)$ augmentation strategy
Randomly initialise student model $S_{\theta}$ and teacher head $h_{\phi}$, or initialise using previous weights if fine-tuning;
$i=0$;
while $i<N_{E}$ do
  Randomly split $X$ into $B$ mini-batches;
  for $x_{b}\in\{X_{1},...,X_{b},...,X_{B}\}$ do
   Generate augmented views: $\mathbf{X}=\text{Aug}(x_{b})$;
   Compute dimensionality reduction objective (Eq.˜12):
   $\mathcal{L}_{\text{dim-red}}(\mathbf{X};\phi)=L_{\text{KL}}(h_{\phi}(T^{c}(\mathbf{X})),T^{c}(\mathbf{X}))+\frac{1}{|\mathbf{X}|}\sum_{i}L_{\text{KL}}(h_{\phi}(T(x_{i})),T(x_{i}))$
   Compute student objective, while freezing $\phi$ (Eq.˜15):
   $\mathcal{L}_{\text{student}}(\mathbf{X};\theta)=L_{\text{cosine}}(S_{\theta}^{c}(\mathbf{X}),h_{\phi_{\text{frozen}}}(T^{c}(\mathbf{X})))+L_{\text{cosine}}(S_{\theta}(\mathbf{X}),h_{\phi_{\text{frozen}}}(T(\mathbf{X})))$
   Combine losses: $\mathcal{L}=\mathcal{L}_{\text{student}}(\mathbf{X};\theta)+\mathcal{L}_{\text{dim-red}}(\mathbf{X};\phi)$;
   Minimise loss $\mathcal{L}$ by updating parameters of $\theta$ and $\phi$;
  end for
  $i=i+1$;
end while
@@ Appendix B Further Results
Method Arch Teacher kNN Linear Backbone Student head Teacher head Teacher Proteus ViT-Ti/14 DINOv2 ViT-S/14 73.1 73.5 79.0 76.1 CosPress ViT-Ti/14 DINOv2 ViT-S/14 74.3 78.8 79.0 76.8 Proteus ViT-Ti/14 DINOv2 ViT-B/14 73.4 73.8 82.1 76.9 CosPress ViT-Ti/14 DINOv2 ViT-B/14 75.6 81.9 82.1 77.9
@@
We observe that training with larger teachers in comparison to the student requires longer training runs. To achieve best results efficiently, the pretrained models from a smaller teacher are used as a starting point. Then, the student and teacher heads are first trained with a frozen pretrained student model (allowing the CosPress teacher heads to minimize the student loss) for 30 epochs. Finally, the models are trained for 300 epochs using the same distillation approach as previously. Table˜S14 shows that this improves the performance of the student models, with CosPress seeing larger improvements in accuracy on ImageNet-1K in comparison to Proteus.
However, this results in poorer results in dense image tasks, like semantic segmentation. Table˜S15 shows that the students trained with a larger teacher network have poorer mIoU for a linear probe on the Pascal VOC 2012 dataset. Undertaking longer distillation runs, or continuing training using a larger image resolution potentially might be helpful in these cases.
@@
Method Arch Teacher mIoU Proteus ViT-Ti/14 DINOv2 ViT-S/14 70.5 CosPress ViT-Ti/14 DINOv2 ViT-S/14 71.1 Proteus ViT-Ti/14 DINOv2 ViT-B/14 68.1 CosPress ViT-Ti/14 DINOv2 ViT-B/14 69.7
@@
The OOD detection results for all of the OpenOOD datasets (yang2022openood) are presented in Table˜S16, Table˜S17 and Table˜S18. The tables report the ROC-AUC for detecting OOD images and the False Positive Rate (FPR) using a 95% threshold for including all in-distribution images. To produce these results, the KNN+ (sun2022outofdistributiondetectiondeepnearest) OOD metric is used to measure the performance of the model backbones. For ImageNet-1K, we sample 1% of the dataset (12,812 images) and set $k=10$ to measure distance to OOD samples. For CIFAR-10/100, we sample 100% of the dataset (50,000 images) and set $k=1$.
@@
Table S17: Specialist models — near OOD detection. Comparison of performance on the OpenOOD benchmark (yang2022openood). The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
@@
Table S18: Specialist models — far OOD detection. Comparison of performance on the OpenOOD benchmark (yang2022openood). The $\uparrow$ means larger values are better and the $\downarrow$ means smaller values are better.
@@ Appendix C Ablation Studies
Method Arch Teacher $\gamma$ kNN CosPress ViT-Ti/14 DINOv2 ViT-S/14 - 74.3 CosPress ViT-Ti/14 DINOv2 ViT-S/14 10 74.3 CosPress ViT-Ti/14 DINOv2 ViT-S/14 100 74.2
@@
We consider the performance of weighting the dimensionality reduction loss in Eq.˜6, rather than freezing the gradients of $\phi$ in the student loss. This gives an alternative loss function
@@
where $\gamma$ is a weighting factor prioritises the dimensionality reduction loss when it is set to be greater than one. Table˜S19 shows that the approach of weighting or freezing gradients leads to similar results.
@@
Method Arch Teacher Loss Metric kNN CosPress ViT-Ti/14 DINOv2 ViT-S/14 Cosine distance 74.3 CosPress ViT-Ti/14 DINOv2 ViT-S/14 MSE 74.2
@@
Table˜S20 considers the impact on performance of using different metrics for the student loss $\mathcal{L}_{\text{student}}$ for Eq.˜15. It is found that using a cosine distance loss leads to slightly better performance in comparison to a mean squared error loss.
@@
Method Arch Teacher $\bm{\tau}$ kNN CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\bm{\left[0.01,0.02,0.03,0.04,0.05,0.06,0.07,0.08,0.09,0.10\right]}$ 74.3 CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\left[0.01\right]$ 74.1 CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\left[0.10\right]$ 74.5 CosPress ViT-Ti/14 DINOv2 ViT-S/14 $\left[0.01,0.10\right]$ 74.3
@@
Table˜S20 shows the performance impact of different sets of temperatures $\bm{\tau}$ in the dimensionality reduction loss for Eq.˜11. It is found that these values have a small impact on performance, with the best set being $\bm{\tau}=\left[0.10\right]$, which obtains slightly better performance that the set of parameters chosen for the paper.